B] <P[AP[B], and of two events A and B Give an example of two events such that P[ A with P[An B] > P[A]P[B].
Answer #1

Consider can the event of rolling a fair dice: cased: - A Getting an even no y B(A)= 1/2 B: . • odd « P(B) = 1/2 but PLANB)=0